OpenStudy (anonymous):

abcd is a rectangle . find the length od the diagona. ac=2x-3 bd=x+5

OpenStudy (anonymous):

Is it 0D diagonal or Of the diagonal?

OpenStudy (anonymous):

of the diagonal

OpenStudy (anonymous):

diagonals of a rectangle are equal ==> ac = bd 2x-3 = x+5 2x - x = 5 + 3 x = 8 so ac = 2(8) - 3 = 16 - 3 = 13 bd = x+5 = 8 + 5 = 13

OpenStudy (anonymous):

that is wronq

OpenStudy (anonymous):

Sqrt[(2x-3)^2+(x+5)^2]

OpenStudy (anonymous):

then you have put the question wrong make the rectangle and you will see that ac and bd are the diagonals.......

OpenStudy (anonymous):

while naming the rectangle, you have to name either clockwise or anticlockwise.. so and c will be always in opposite corners and the line joining them will be the diagonal.... similarly for ac....

OpenStudy (anonymous):

see attached file...